Circular No. 6316

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

HD 3346 R. Noyes, S. Korzennik, P. Nisenson, S. Jha, and M. Krockenberger,

Smithsonian Astrophysical Observatory; T. Brown, T. Kennelly, and C. Rowland, High Altitude Observatory; and S. Horner, Pennsylvania State University, report the detection of large short-term radial-velocity variations in the K5 III star HD 3346 = HR 152. Irregularly spaced observations were made with the Advanced Fiber Optic Echelle (AFOE) spectrometer at the 1.5-m telescope of the Whipple Observatory during the last quarters of 1993, 1994, and 1995, with more intensive observations in December 1995 and February 1996. A long-period radial-velocity variation similar to that already reported by McClure et al. (1985, PASP 97, 740) is seen with amplitude of order 500 m/s and period of order 650 days. This is consistent with a companion to HD 3346 at orbital distance ∼ 2.5 AU. If the mass of HD 3346 is 5 M_{⊙}, then this already-reported companion would have projected mass m sin i ∼ 60 Jupiter masses. The AFOE data reveal the presence of shorter-term radial-velocity variations superimposed on the 650-day variation. The amplitude of the variations is in the range 150–300 m/s. Near-nightly observations in December 1995 and February 1996 show the variations to be consistent with those produced by a second orbital companion. The window function of the data would allow for its period to be near 14, 18, 24, or 40 days. Such a short period would imply that the second companion’s orbit would have a semimajor axis between 0.2 and 0.4 AU; the amplitude of the variation implies that (again, if HD 3346 has mass 5 M_{⊙}) this second companion has m sin i ∼ 10 Jupiter masses. Acoustic pulsations may be an alternative source of the short-term radial-velocity variations. However, the fundamental period of acoustic pulsations in giants is expected to be only a few days; the present data could be explained by acoustic pulsations only if the star has a mass much lower than suggested by standard stellar-evolution theory. Spurious signals due to rotation of starspots appear to be unlikely because of the low reported rotational velocity for HD 3346. A definitive period for the short-period radial-velocity variation may be determinable if other contemporaneous precise radial-velocity observations of this star exist or can be obtained before the star disappears behind the sun for this observing season.

1996 February 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green